# Collimation of Fast Radio Burster 20201124A; Repeaters vs. Apparent Non-Repeaters

### v1

@@ Abstract
The recent report of a period in the active repeating Fast Radio Burster 20201124A and measurement of its spindown rate place bounds on the solid angle of its emission on the basis of energetics. The bound depends on the (unknown) efficiency of conversion of rotational energy to coherent radio emission and implies a lower bound on the Lorentz factor of the radiating charges. Bursts may be emitted along the magnetic dipole axis, in repeaters aligned with the rotational axis and the line of sight but misaligned in apparent non-repeaters. This may explain the difficulty of finding periodicity in repeaters and the low duty cycle of apparent non-repeaters.
@@ 1 Introduction
Fast Radio Burst (FRB) energetics remain enigmatic. The report (, 2025) of a 1.7 s period during two observations, 37 days apart, of the very active repeating FRB 20201124A implied that at least this FRB is produced by a rotating magnetic neutron star, and the similarity of the two reported periods supported their reality. The statistical significance of these detections was recently challenged (Gazith & Zackay, 2025). Zhang et al. (2025) reported a correlation between the the times of the claimed periodicities and changes in Rotation Measure, which may support the reality of the periodicities. The present paper, a preliminary version of which was published (Katz, 2025) before that challenge, discusses inferences that may be drawn if such periodicity is confirmed, either for FRB 20201124A or for some other repeating FRB.
@@ 4 Discussion
Comparing the spindown rate and flux of FRB 20201124A estimates the collimation of its radiation and bounds the Lorentz factor of its emitting charges. It also suggests that the radiation is emitted along open field lines because their large radii of curvature imply a comparatively large volume can contribute to collimated radiation, relaxing the constraint on the magnetic field that confines the radiating charges. Alignment of the magnetic dipole and rotation axes of repeating FRB would explain the difficulty of finding their rotational periodicity because radiation from a field structure azimuthally symmetric around the rotation axis would not be rotationally modulated. It is implausible that neutron star fields are accurately dipolar near the surface, so this argument suggests that FRB are emitted at many neutron star radii. Emission along open field lines is consistent with, but does not require, that the periodic activity of some repeating FRB results from free precession of the neutron star (Zanazzi & Lai, 2020).
FRB emission along the magnetic dipole axis may also explain the distinction between repeating and non-repeating FRB: If aligned magnetic and rotational axes point in our direction, we may observe all the source’s bursts as a repeating FRB with intrinsic duty factor DF. If the magnetic and rotational axes are misaligned we can only observe bursts when, fortuitously and infrequently, the magnetic axis points toward us. We can observe only a fraction ${\cal O}(\Omega/4\pi)$ of all active repeaters close and powerful enough for detection because the others are beamed away from us. We can observe a large fraction of sufficiently close and powerful mis-aligned sources, but only as apparent non-repeaters with duty factors ${\cal O}(\text{DF}\Omega/4\pi)$; approximating the magnetic axes of the apparent non-repeaters as uniformly distributed on the sky, $\Omega/4\pi$ is the fraction of the time they are pointed in our direction. This is consistent with constraints (Katz, 2024b) on their duty factors, but both $\Omega$ and DF are uncertain even to order of magnitude.

### v2

@@ Abstract
The recent report of a period in the active repeating Fast Radio Burster 20201124A and of its spindown rate place bounds on the solid angle of its emission on the basis of energetics. The bound depends on the (unknown) efficiency of conversion of rotational energy to coherent radio emission and implies a lower bound on the Lorentz factor of the radiating charges. Bursts may be emitted along the magnetic dipole axis, in repeaters aligned with the rotational axis and the line of sight but misaligned in apparent non-repeaters. This may explain the difficulty of finding periodicity in repeaters and the low duty cycle of apparent non-repeaters.
@@ 1 Introduction
Fast Radio Burst (FRB) energetics remain enigmatic. The report (, 2025) of a 1.7 s period during two observations, 37 days apart, of the very active repeating FRB 20201124A implied that at least this FRB is produced by a rotating magnetic neutron star, and the similarity of the two reported periods supported their reality. The statistical significance of these detections was recently challenged (Gazith & Zackay, 2025). Zhang et al. (2025) reported a correlation between the the times of the claimed periodicities and changes in Rotation Measure, which may support the reality of the periodicities. The present paper, a preliminary version of which was published (Katz, 2025) before that challenge, discusses inferences that may be drawn if periodicity and spindown are confirmed, either for FRB 20201124A or for some other repeating FRB.
@@ 4 Discussion
Comparing the spindown rate and flux of FRB 20201124A estimates the collimation of its radiation and bounds the Lorentz factor of its emitting charges. It also suggests that the radiation is emitted along open field lines, as previously suggested by Beniamini & Kumar (2025), because their large radii of curvature imply a comparatively large volume can contribute to collimated radiation, relaxing the constraint on the magnetic field that confines the radiating charges.
Alignment of the magnetic dipole and rotation axes of repeating FRB would explain the difficulty of finding their rotational periodicity because radiation from a field structure azimuthally symmetric around the rotation axis would not be rotationally modulated. It is implausible that neutron star fields are accurately dipolar near the surface, so this argument suggests that FRB are emitted at many neutron star radii. Emission along open field lines is consistent with, but does not require, that the periodic activity of some repeating FRB results from free precession of the neutron star (Zanazzi & Lai, 2020).
FRB emission along the magnetic dipole axis may also explain the distinction between repeating and non-repeating FRB: If aligned magnetic and rotational axes point in our direction, we may observe all the source’s bursts as a repeating FRB with intrinsic duty factor DF. If the magnetic and rotational axes are misaligned we can only observe bursts when, fortuitously and infrequently, the magnetic axis points toward us. We can observe only a fraction ${\cal O}(\Omega/4\pi)$ of all active repeaters close and powerful enough for detection because the others are beamed away from us (Beniamini & Kumar, 2025). We can observe a large fraction of sufficiently close and powerful mis-aligned sources, but only as apparent non-repeaters with duty factors ${\cal O}(\text{DF}\Omega/4\pi)$; approximating the magnetic axes of the apparent non-repeaters as uniformly distributed on the sky, $\Omega/4\pi$ is the fraction of the time they are pointed in our direction. This is consistent with constraints (Katz, 2024b) on their duty factors, but both $\Omega$ and DF are uncertain even to order of magnitude.